OpenStudy (anonymous):

Find all real solutions of the equation exactly. x^4-3x^3-10=0

OpenStudy (anonymous):

Factor: \[(x-3)x ^{3}= 10\] 2 real and 2 imaginary solutions. Real: \[x \approx -1.32257, 3.28269\]

OpenStudy (anonymous):

Are you sure it's \(x^4-3x^3-10=0\), because such a problem should be solved numerically, not analytically. It would too difficult to find exact solutions.

OpenStudy (anonymous):

The problem is probably \(x^4-3x^2-10=0\), which can be easily solved by letting \(u=x^2\), and then use the quadratic formula to solve it. Don't forget to substitute back for x.